how many grams of H3BO3 (s) will be required for the complete reaction of 1.55 ml of 28.6 M HF (aq)?
1
Expert's answer
2019-10-24T08:00:34-0400

"H_3BO_3 + HF = H[BF_4] + H_2O"

As per reaction ,

"M_2V_2=" Moles of "HF=" Moles of "H_3BO_3=28.6\times 1.55\ milimoles=" "44.33\times 10^{-3} \ moles"

Weight of "H_3BO_3= Moles\times Molar \ mass="

"44.33 \times 61.83 \times 10^{-3}=" "2.741 \ gm"